# GVE-Leiden: Fast Leiden Algorithm for Community Detection in Shared Memory Setting

### 21 Dec 2023 v1

@@ 5.2. Comparing Performance of GVE-Leiden
Figure 4(a) shows the runtimes of the original Leiden, igraph Leiden, NetworKit Leiden, and GVE-Leiden on each graph in the dataset. On the sk-2005 graph, GVE-Leiden finds communities in $10.8$ seconds, and thus achieve a processing rate of $352$ million edges/s. Figure 4(b) shows the speedup of GVE-Leiden with respect to each implementation mentioned above. GVE-Leiden is on average 373×373$\times$373 ×, 86×86$\times$86 ×, and 7.2×7.2$\times$7.2 × faster than the original Leiden, igraph Leiden, and NetworKit Leiden respectively. Figure 4(c) shows the modularity of communities obtained with each implementation. GVE-Leiden on average obtains $0.1\%$ lower modularity than the original Leiden and igraph Leiden, and $26\%$ higher modularity than NetworKit Leiden (especially on road networks and protein k-mer graphs). Finally, Figure 4(d) shows the fraction of disconnected communities obtained with each implementation. Here, the absence of bars indicates the absence of disconnected communities. Communities identified by GVE-Leiden on average have $\times8$, 145×145$\times$145 ×, and 0.76×0.76$\times$0.76 × disconnected communities than the original Leiden, igraph Leiden, and NetworKit Leiden respectively. While this compares unfavorably with the original Leiden and igraph Leiden (especially on social networks, road networks, and protein k-mer graphs), it may be simpler to split the disconnected communities obtained from GVE-Leiden as a post-processing step. We would like to address this issue some time in the future.
Next, we compare the performance of GVE-Leiden with GVE-Louvain (Sahu, 2023), our optimized parallel implementation of the Louvain method. As above, for each graphs in the dataset, we run both algorithms 5 times to minimize measurement noise, and report the averages in Figures 5(a), 5(b), 5(c), and 5(d). Figure 5(a) shows the runtimes of GVE-Louvain and GVE-Leiden on each graph in the dataset. Figure 5(b) shows the speedup of GVE-Leiden with respect to GVE-Louvain. GVE-Leiden is on average $36\%$ slower than GVE-Louvain. This increase in computation time is a trade-off for obtaining significantly fewer disconnected communities, as given below. Figure 5(c) shows the modularity of communities obtained with GVE-Louvain and GVE-Leiden. GVE-Leiden on average obtains $0.3\%$ higher modularity than GVE-Louvain. Finally, Figure 5(d) shows the fraction of disconnected communities obtained with GVE-Louvain and GVE-Leiden. Communities identified by GVE-Leiden on average have an $11$-fold decrease in the number of disconnected communities than GVE-Louvain.
@@ 5.3. Analyzing Performance of GVE-Leiden
We now analyze the performance of GVE-Leiden. The phase-wise and pass-wise split of GVE-Leiden is shown in Figures 6(a) and 6(b) respectively. Figure 6(a) indicates that GVE-Leiden spends most of its runtime in the local-moving and refinement phases on web graphs, road networks, and protein k-mer graphs, while it devotes majority of its runtime in the aggregation phase on social networks. The pass-wise split, shown in Figure 6(b), indicates that the first pass dominates runtime on high-degree graphs (web graphs and social networks), while subsequent passes prevail in execution time on low-degree graphs (road networks and protein k-mer graphs).
On average, $37\%$ of GVE-Leiden’s runtime is spent in the local-moving phase, $27\%$ in the refinement phase, $24\%$ is spent in the aggregation phase, and $12\%$ is spent in other steps (initialization, renumbering communities, looking up dendrogram, and resetting communities) of the algorithm. Further, $73\%$ of the runtime is spent in the first pass of the algorithm, which is the most expensive pass due to the size of the original graph (later passes work on super-vertex graphs). We also observe that graphs with lower average degree (road networks and protein k-mer graphs) and graphs with poor community structure (such as com-LiveJournal and com-Orkut) have a larger $\text{runtime}/|E|$ factor, as shown in Figure 7.
@@ 5.4. Strong Scaling of GVE-Leiden
Finally, we measure the strong scaling performance of GVE-Leiden. To this end, we adjust the number of threads from $1$ to $64$ in multiples of $2$ for each input graph, and measure the overall time taken for finding communities with GVE-Leiden, as well as its phase splits (local-moving, refinement, aggregation, others), five times for averaging. The results are shown in Figure 8. With 32 threads, GVE-Leiden obtains an average speedup of 10.9×10.9$\times$10.9 × compared to running with a single thread, i.e., its performance increases by 1.6×1.6$\times$1.6 × for every doubling of threads. Scaling is limited due to the various sequential steps/phases in the algorithm. At 64 threads, GVE-Leiden is impacted by NUMA effects, and offers speedup of only 12.9×12.9$\times$12.9 ×.
@@ 6. Conclusion
On a system equipped with two 16-core Intel Xeon Gold 6226R processors, our implementation of the Leiden algorithm, referred to as GVE-Leiden, attains a processing rate of $352M$ edges per second on a $3.8B$ edge graph. It surpasses the original Leiden implementation, igraph Leiden, and NetworKit Leiden by factors of 373×373$\times$373 ×, 86×86$\times$86 ×, and 7.2×7.2$\times$7.2 × respectively. GVE-Leiden identifies communities of equivalent quality to the first two implementations, and $26\%$ higher quality than NetworKit. Doubling the number of threads results in an average performance scaling of 1.6×1.6$\times$1.6 × for GVE-Leiden. In comparison to GVE-Louvain (Sahu, 2023), our optimized parallel Louvain implementation, GVE-Leiden reduces internally-disconnected communities by a factor of $11$ with only a $36\%$ increase in runtime.
However, on average, communities identified by GVE-Leiden exhibit $\times8$, 145×145$\times$145 ×, and 0.76×0.76$\times$0.76 × more disconnected communities compared to the original Leiden, igraph Leiden, and NetworKit Leiden, respectively. This disparity is particularly noticeable on social networks, road networks, and protein k-mer graphs. Despite this drawback, a potential solution is to address the issue by post-processing, where disconnected communities obtained from GVE-Leiden can be split. We acknowledge this concern, and plan to explore solutions in future work.
@@ Acknowledgements.
I would like to thank Prof. Kishore Kothapalli and Prof. Dip Sankar Banerjee for their support.

### 12 Mar 2024 v4

@@ 5.2. Comparing Performance of GVE-Leiden
Figure 4(a) shows the runtimes of the original Leiden, igraph Leiden, NetworKit Leiden, and GVE-Leiden on each graph in the dataset. On the sk-2005 graph, GVE-Leiden finds communities in $9.4$ seconds, and thus achieve a processing rate of $403$ million edges/s. Figure 4(b) shows the speedup of GVE-Leiden with respect to each implementation mentioned above. GVE-Leiden is on average 436×436$\times$436 ×, 104×104$\times$104 ×, and 8.2×8.2$\times$8.2 × faster than the original Leiden, igraph Leiden, and NetworKit Leiden respectively. Figure 4(c) shows the modularity of communities obtained with each implementation. GVE-Leiden on average obtains $0.3\%$ lower modularity than the original Leiden and igraph Leiden, and $25\%$ higher modularity than NetworKit Leiden (especially on road networks and protein k-mer graphs). Finally, Figure 4(d) shows the fraction of disconnected communities obtained with each implementation. Here, the absence of bars indicates the absence of disconnected communities. Communities identified by the original Leiden, igraph Leiden, and NetworKit Leiden have on average $1.3\times 10^{-4}$, $7.9\times 10^{-5}$, and $1.5\times 10^{-2}$ fraction of disconnected communities, while none of the communities identified by GVE-Leiden are internally-disconnected. As the Leiden algorithm guarantees the absence of disconnected communities (Traag et al., 2019), those observed with the original Leiden, igraph Leiden, and NetworKit Leiden are likely due to implementation issues.
Next, we compare the performance of GVE-Leiden with GVE-Louvain (Sahu, 2023), our parallel implementation of the Louvain method. As above, for each graph in the dataset, we run both algorithms 5 times to minimize measurement noise. Figure 5(a) shows the runtimes of GVE-Louvain and GVE-Leiden on each graph in the dataset. Figure 5(b) shows the speedup of GVE-Leiden with respect to GVE-Louvain. GVE-Leiden is on average $13\%$ slower than GVE-Louvain. This increase in computation time is a trade-off for identifying communities that are not internally-disconnected, as given below. Figure 5(c) shows the modularity of communities obtained with GVE-Louvain and GVE-Leiden. GVE-Leiden on average obtains the same modularity as GVE-Louvain. Finally, Figure 5(d) shows the fraction of internally-disconnected communities obtained with GVE-Louvain and GVE-Leiden. Communities identified by GVE-Louvain on average have $4.0\%$ disconnected communities, while GVE-Leiden has none.
@@ 5.3. Analyzing Performance of GVE-Leiden
We now analyze the performance of GVE-Leiden. The phase-wise and pass-wise split of GVE-Leiden is shown in Figures 6(a) and 6(b) respectively. Figure 6(a) reveals that GVE-Leiden devotes a significant portion of its runtime to the local-moving and refinement phases on web graphs, road networks, and protein k-mer graphs, while it dedicates majority of its runtime in the aggregation phase on social networks. The pass-wise split, shown in Figure 6(b), indicates that the first pass is time-intensive for high-degree graphs (web graphs and social networks), while subsequent passes take precedence in execution time on low-degree graphs (road networks and protein k-mer graphs).
On average, GVE-Leiden spends $46\%$ of its runtime in the local-moving phase, $19\%$ in the refinement phase, $20\%$ in the aggregation phase, and $15\%$ in other steps (initialization, renumbering communities, dendrogram lookup, and resetting communities). Further, $63\%$ of the runtime is consumed by the first pass of the algorithm, which is computationally demanding due to the size of the original graph (subsequent passes operate on super-vertex graphs). We also observe that graphs with lower average degree (road networks and protein k-mer graphs) and those with poor community structure (e.g., com-LiveJournal and com-Orkut) exhibit a higher $\text{runtime}/|E|$ factor, as shown in Figure 7.
@@ 5.4. Strong Scaling of GVE-Leiden
Finally, we assess the strong scaling performance of GVE-Leiden. In this analysis, we vary the number of threads from $1$ to $64$ in multiples of $2$ for each input graph, and measure the total time taken for GVE-Leiden to identify communities, encompassing its phase splits (local-moving, refinement, aggregation, and others), repeated five times for averaging. The results are shown in Figure 8. With 32 threads, GVE-Leiden achieves an average speedup of 11.4×11.4$\times$11.4 × compared to a single-threaded execution, indicating a performance increase of 1.6×1.6$\times$1.6 × for every doubling of threads. Nevertheless, scalability is restricted due to the sequential nature of steps/phases in the algorithm. At 64 threads, GVE-Leiden is affected by NUMA effects, resulting in a speedup of only 16.0×16.0$\times$16.0 ×.
@@ 6. Conclusion
On a system equipped with two 16-core Intel Xeon Gold 6226R processors, our implementation of the Leiden algorithm, referred to as GVE-Leiden, attains a processing rate of $403M$ edges per second on a $3.8B$ edge graph. It surpasses the original Leiden implementation, igraph Leiden, and NetworKit Leiden by factors of 436×436$\times$436 ×, 104×104$\times$104 ×, and 8.2×8.2$\times$8.2 × respectively. GVE-Leiden identifies communities of equivalent quality to the first two implementations, and $25\%$ higher quality than NetworKit. Doubling the number of threads results in an average performance scaling of 1.6×1.6$\times$1.6 × for GVE-Leiden. In comparison to GVE-Louvain (our parallel Louvain implementation) (Sahu, 2023), the original Leiden, igraph Leiden, and NetworKit Leiden, GVE-Leiden completely eliminates internally-disconnected communities.
In the previous version of this report, we implemented the refinement phase of the Leiden algorithm utilizing a constrained move procedure, which does not guarantee the absence of disconnected communities. In this current version of the report, we have transitioned to employing the constrained merge procedure alongside atomics to ensure no internally-disconnected communities.
@@ Acknowledgements.
I would like to thank Prof. Kishore Kothapalli, Prof. Dip Sankar Banerjee, Vincent Traag, Geerten Verweij, and Fabian Nguyen for their support.
Appendix A Appendix
A.1. Community labels of super-vertices
We also attempt two different variations of Parallel Leiden algorithm, one where the community labels of super-vertices (upon aggregation) is based on the local-moving phase (move-based), and the other where the community labels of super-vertices is based on the refinement phase (refine-based). Our observations indicate that both approaches have roughly the same runtime and modularity on average, as indicated by Figures 9 and 10. Accordingly, we stick to the move-based approach, which is the one recommended by Traag et al. (Traag et al., 2019). However, refine-based approach may be more suitable for the design of dynamic Leiden algorithm (for dynamic graphs).
Figure 9. Average relative runtime for move-based and refine-based communities for super-vertices upon aggregation with parallel Leiden algorithm, for all graphs in the dataset.
Figure 10. Average modularity for move-based and refine-based communities for super-vertices upon aggregation with parallel Leiden algorithm, for all graphs in the dataset.